The sound level of a 1 W/m squared intensity is (in dB)
A. 1
B. 10
C .120
D. 130
E. 200
Answer #1

here,

Intensity level, in dB = 10 log(I/Io) where log = log base 10

I0 = 1 * 10^-12 W/m^2

so,

the sound level in dB = 10 * log(1 /(1 * 10^-12))

dB = 120 dB

the sound level in dB is C) 120